40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 240 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  4909   Wed Jun 29 11:45:54 2011 steveUpdateGeneralwhere are we storing PDs ?

Photo diodes stored in the east arm cabinet E4:  ALL PDs  meet here, fast  or slow......

Attachment 1: P1070921.JPG
P1070921.JPG
  4908   Wed Jun 29 11:25:07 2011 NicoleSummarySUSWeekly Summary of Work

Update of Week 3 Work:

-I've finished reading The Art of Electronics Ch 1, 2, and 4.

-The mechanical stage for the horizontal displacement measurements is set up.

-I've opened up the circuit box for the quad photodiode and am currently working on the circuit diagram for the box and for the quad photodiode sensors.

 

Later this week, I plan to finish the circuit diagrams and figure out how the circuits work with the four inputs. I also plan to start working on my first

progress report.

 

  4907   Wed Jun 29 11:13:13 2011 JenneUpdateSUSissues in the current quad maglev system

magnetQuote


(1) DC magnetic force is imbalanced, and we found that one pair has a stronger DC force than others. This should
be able to solved simply by replacing them with magnets have comparable strength to others.

 I don't know if this would work, but it might be worth a try:

You've achieved single levitation before, with fairly good stability.  Can you try taking each magnet + coil and finding the DC coil current required to hold a mass at a given position?  If you can hold the same mass at the same place with all the different magnets+coils, then you're exerting the same force against gravity, so your DC forces are balanced. 

  4906   Wed Jun 29 01:23:21 2011 haixingUpdateSUSissues in the current quad maglev system

Here I show several issues that we have encountered in the quad magnetic levitation system. It would be great if you can give
some suggestions and comments (Poor haixing is crying for help)

The current setup is shown by the figure below (I took the photo this morning):

config_2.png

Basically, we have one heavy load which is rigidly connected to a plane that we try to levitate. On corners of the
plane, there are four push-fit permanent magnets. Those magnets are attracted by four other magnets which are
mounted on the four control coils (the DC force is to counteract the DC gravity). By sensing the position of the plane
with four OSEMs (there are four flags attached on the plane), we try to apply feedback control and levitate the plane.
We have made an analog circuit to realize the feedback, but it is not successful. There are the following main issues
that need to be solved:

(1) DC magnetic force is imbalanced, and we found that one pair has a stronger DC force than others. This should
be able to solved simply by replacing them with magnets have comparable strength to others.

(2) The OSEM not only senses the vertical motion, but also the translational motion. One possible fast solution is to
cover the photodiode and only leave a very thin vertical slit so that a small translational motion is not sensed.
Maybe this is too crappy. If you have better ideas, please let me know. Koji suggested to use reflective sensing
instead of OSEM, which can also solve the issue that flags sometimes touche the hole edge of the OSEM and
screw up the sensing.

(3) Cross coupling among different degrees of freedom. Basically, even if the OSEM only senses the vertical motion,
the motion of four flags, which are rigidly connected to the plane, are not independent. In the ideal case, we only
need to control pith, yaw and vertical motion, which only has three degrees of freedom, while we have four sensing outputs
from four OSEMs. This means that we need to work out the right control matrix. Right now, we are in some kind of dilemma.
In order to obtain the control matrix, we first have to get the sensing matrix or calibrate the cross coupling; however, this is
impossible if the system is unstable. This is very different from the case of quad suspension control used in LIGO,
in which the test mass is stable suspended and it is relatively easy to measure the cross coupling by driving the test mass
with coils. Rana suggested to include a mechanical spring between the fixed plane and levitated plane, so that
we can have a stable system to start with. I tried this method today, but I did not figure out a nice way to place the spring,
as we got a hole right in the middle of the fixed plane to let the coil connectors go though. As a first trial, I plan to
replace the stop rubber band (to prevent the plane from getting stuck onto the magnets) shown in the figure with mechanical
springs. In this case, the levitated plane is held by four springs instead of one. This is not as good as one, because
of imbalance among the four, but we can use this setup, at least, to calibrate the cross coupling. Let me know if you come
up better solution.

After those issues are solved, we can then implement Jamie's Cymac digital control, which is now under construction,
to achieve levitation.

  4905   Wed Jun 29 00:35:36 2011 KojiUpdateLSCnew LSC overview screen 80% done

New LSC screen is 80% completed.

It is now accessible from the LSC menu of "sitemap".

Most of the part in the screen is clickable such that it launches another screen depending on the location of the click.

 

The bottom part of the screen still need some work.

RFPD screen is temporary

LSC control screen is also temporary

DAC overflow indicators are still broken.

Channel assignment of the whitening filters are arbitrary so far.

 

Attachment 1: LSC_OVERVIEW.png
LSC_OVERVIEW.png
  4904   Tue Jun 28 22:36:04 2011 JamieUpdateSUSChecking binary switching of SUS whitening filter

I have been checking the binary output switching for the SUS whitening filters. It appears that the whitening switching is working for (almost) all the vertex suspensions (BS, ITMX, ITMY, PRM, SRM), but not for the ETMs.

The table below lists the output from my switch-checking script (attached). The script uses the SUS digital lockin to drive one coil and measure the same coil's OSEM response, repeating for each coil/OSEM pair. I used a lockin drive frequency of about 10 Hz, at which the whitening filter should have 10 db of gain.

All but one of the vertex OSEMS show the proper response (~10db gain at 10Hz) when the whitening is switched on from the digital controls. ITMY UL appears to not be switching, which I fear is due to my electronics fail noted in my previous log post.  The ETMs are clearly not switching at all.

I will try to get the ETM switching working tomorrow, as well as try to asses what can be done about the ITMY UL switch.  After that I will work on confirming the coil drive dewhite switching.

lockin settings

freq: 10.123 Hz
amp: 10000
I/Q filters: 0.1 Hz LP, 4-pole butterworth

response

BS
ul : 3.31084503062 = 10.3987770676 db
ll : 3.34162124753 = 10.4791444741 db
sd : 3.43226254574 = 10.7116100229 db
lr : 3.28602651913 = 10.3334212798 db
ur : 3.29361593249 = 10.3534590969 db

ITMX
ul : 3.37499773336 = 10.5654697099 db
ll : 3.2760924572  = 10.3071229966 db
sd : 3.13374799272 =  9.9212813757 db
lr : 3.28133776018 = 10.3210187243 db
ur : 3.37250879937 = 10.5590618297 db

ITMY
ul : 0.99486434364 = -0.0447226830807 db
ll : 3.39420873724 = 10.6147709414 db
sd : 3.88698713176 = 11.7922620572 db
lr : 3.357123865   = 10.5193473069 db
ur : 3.37876008179 = 10.5751470918 db

PRM
ul : 3.26758918055 = 10.2845489876 db
ll : 3.32023820566 = 10.4233848529 db
sd : 3.25205538857 = 10.2431586766 db
lr : 3.24610681962 = 10.227256141  db
ur : 3.31311970305 = 10.4047425446 db

SRM
ul : 3.30506423619 = 10.3835980943 db
ll : 3.28152094133 = 10.3215036019 db
sd : 3.08566647696 =  9.7869796462 db
lr : 3.30298270419 = 10.378125991  db
ur : 3.3012249406  = 10.3735023505 db

ETMX
ul : 0.99903400106 = -0.00839461539757 db
ll : 0.99849991349 = -0.0130393683795 db
sd : 1.00314092883 =  0.0272390056874 db
lr : 1.00046493718 =  0.00403745453682 db
ur : 1.00265600785 =  0.0230392084558 db

ETMY
ul : 1.00223179107 =  0.0193634913327 db
ll : 0.96755532811 = -0.286483823189 db
sd : 1.00861855271 =  0.0745390477589 db
lr : 1.05718545676 =  0.483023602007 db
ur : 0.99777406174 = -0.0193558045143 db
Attachment 1: botest.py
#!/usr/bin/env python

import sys
import os
import subprocess
import time
import pickle
from numpy import *
import nds
import matplotlib
... 207 more lines ...
  4903   Tue Jun 28 22:07:46 2011 JamieHowToSAFETYEurocrate extender card fried (ie. Jamie did a very bad thing)

IMPORTANT ELECTRONICS SAFETY LESSON TO FOLLOW

Yesterday, I fried the +15 V power supply rail on one of the Eurocrate extender cards while I was checking out the binary switching in the 1X5 rack.  I will describe what I did it in the hopes that everyone else will be less stupid than me.

I wanted to monitor the voltage across a resistor on the suspension OSEM whitening board.  Since I knew that both sides of the resistor would be at non-zero voltage (including possibly at the power-supply rail), I used a battery-operated scope with floating inputs, so that the scope would not try to pull the probe shield to ground.  That should be OK, although not recommended, as you'll see, because you must be very careful to make sure that the scopes inputs are indeed floating.

Let's call the original signal 'A'.  The trouble came when I then connected another signal (B), whose shield was connected to the ground on the whitening board, to the scope.  Apparently the grounds on the scope inputs are connected, or were in the configuration I was using.  When I connected the signal B, B's ground shorted A's shield to ground, which had been sitting at the +15V rail.  That short circuit then fried the +15V supply line on the extender card I was using (escaping magic smoke was detected).  Thankfully this only blew the extender card, and not the Eurocrate or the Eurocrate power supply or the whitening board or the scope etc, all of which would have been much worse.

The moral of the story is to be very careful when connecting power supply voltages to the shield or ground of a scope.  In short, don't do it.  I didn't ultimately need to, since I could have found other ways to measure the same signal.

 

  4902   Tue Jun 28 21:05:05 2011 JamieUpdateSUSSUS control model updated

I have updated the sus_single_control model, adjusting/cleaning up/simplifying the LSC/POS input signals, and routing new signals to the lockins. Notably one of POS inputs to the part ("lockin_in") was eliminated (see below).

The 6 inputs to the TO_COIL output matrix are now:

LSCPOS + OFFSET + ALT_POS_IN
ASCPIT + OFFSET + SUSPIT + OLPIT
ASCYAW +OFFSET + SUSYAW + OLYAW
SIDE
LOCKIN1
LOCKIN2

The ALT_POS input is used only by the ETMs for the green locking. Just outside of the sus_single_control library part in the ETM models are the green locking controls, consisting of the ETM?_ALS filter bank and the ETM?_GLOCKIN lockin, the outputs from which are summed and fed into the aforementioned ALT_POS input.

As for the SUS lockins (LOCKIN1 and LOCKIN2 in the library model), their input matrix now gets the direct inputs from the OSEMS (before filtering) and the outputs to the coils, after all filtering. These will aid in doing binary output switching tests.

All suspension models (c1sus, c1scx, c1scy) have been rebuild and restarted so that they reflect these changes.

  4901   Tue Jun 28 16:52:37 2011 SonaliUpdateGreen LockingRouting of fibre to PSL complete.

1. Suresh and I completed the alignment of the fibre and the three mirrors on the ETMY table.

2. We managed to get an output beam power of around 60% using the Ophir(Orion/PD) power meter to finetune the alignment. The power of the input beam is 74.4 mW and of the output beam is 38.5 mW.

3. The coupler on the output side of the fibre which had been put there to help in the alignment has been removed.

4. The picture of the ETMY layout as of now has been attached.

5. The labels A stands for the mirror used to turn the beam direction and B and C stand for the three mirrors used in the alignment of the beam into the coupler,D.(attachment 3).

6. The fibre we used is 50m in length which was barely sufficient to reach the PSL table.

7. So, the fibre has been routed to the PSL table using the fibre tray running below the Y-arm tube as this was the shortest route possible(even though it is a rather acccident prone zone).

8. The fibre has been tied down at regular intervals so that it does not get snagged and pulled up inadvertently.

9. We will start with the preparation of the layout of the PSL table to superpose the two beams on Monday.

Attachment 1: coupled_fibre.jpg
coupled_fibre.jpg
Attachment 2: the_fibre_route.jpg
the_fibre_route.jpg
Attachment 3: ETMY_aftr_fibre_coupling2.png
ETMY_aftr_fibre_coupling2.png
  4900   Tue Jun 28 15:23:21 2011 steveUpdateElectronicsCoilcraft RF-design kits are restocked

Our design kits are full again. They are waiting for a new brilliant PD design.

Attachment 1: P1070917.JPG
P1070917.JPG
Attachment 2: P1070915.JPG
P1070915.JPG
  4899   Tue Jun 28 15:20:08 2011 kiwamuUpdateLSCsensing matrix measurement in PRMI configuration

Here is the result of the measurement of the sensing matrix in the PRMI configuration.

If we believe the resultant matrix, it is somewhat different from what we expected from a finesse simulation (summary of simulated sensing matrix).

 


(Motivation)

As a part of the DRMI test plan, we wanted to check the sensing matrices and consequently diagonalize the LSC input matrix.

The matrix of the DRMI configuration has been measured (#4857), but it was a bit too complicated as a start point.

So first in order to make sure we are doing a right measurement, we moved onto a simpler configuration, that is PRMI.

 

(measurement)

The technique I used was the same as before (#4857) except for the fact that SRM wasn't included this time.

   - PRC was locked to the carrier resonant point. The UGF of MICH and PRC were ~ 110 Hz and 200 Hz respectively.

   - Longitudinally shook BS, ITMs and PRM at 283.103 Hz with an amplitude of 1000 counts using the LOCKIN oscillator in C1LSC.

   - Took the I and Q phase signals from the LOCKIN outputs.

The table below is the raw data obtained from this measurement :

rawmatrix.png

 

(Conversion of matrix)

 With the matrix shown above, we should be able to obtain the sensing matrix which gives the relation between displacements in each DOF to each signal port.

The measured matrix connects two vectors, that is,

       (signal port vector) = [Measured raw matrix] (SUS actuation vector),   -- eq.(1)

where

       (signal port vector) = (AS55_I, AS55_Q, REFL11_I, REFL11_Q)T   in unit of [counts],

    (SUS actuation vector) = (BS, ITMX, ITMY, PRM)T   in units of [counts].

Now we break the SUS actuation vector into two components,

       (SUS actuation vector [counts])  = (actuator response matrix [m/counts])-1 * (MICH, PRM [m] )^T   -- eq.(2)

 where

       (actuator response matrix) =  2.05x10-13 * ( [1   ,  0.217, -0.216,   0  ],

                                                 [ 0.5,  0.109 -0.108, 0.862]  )  in unit of [m/counts]

These values are coming from the actuator calibration measurement.

In the bracket all the values are normalized such that BS has a response of 1 for MICH actuation.

Combining eq.(1) and (2) gives,

     (signal port vector) = (measured raw matrix) * (actuator response matrix)-1 * (MICH, PRM)T

And now we define the sensing matrix by

     (sensing matrix) = (measured raw matrix) * (actuator response matrix)-1

The sensing matrix must be 4x2 matrix.

For convenience I then converted the I and Q signals of each port into the absolute value and phase.

       ABS = sqrt((AAA_I)2 +(AAA_Q)2 ),

       PHASE = atan (AAA_Q / AAA_I),

where AAA is either AS55 or REFL11.

 

(Resultant matrix)

The table below is the resultant sensing matrix.

ABS represents the strength of the signals in unit of [cnts/m], and PHASE represents the demodulation phases in [deg].

sensmat.png

There are several things which I noticed :

   - The demodulation phase of MICH=>AS55 and PRC=>REFL11 are close to 0 or 180 deg as we expected.

      This is a good sign that the measurement is not something crazy.

   - AS55 contains a big contribution from PRC with a separation angle of 152 deg in the demodulation phase.

     In AS55 the signal levels of MICH and PRC were the same order of magnitude but PRC is bigger by a factor of ~4.

     However the finesse simulation (see wiki page) shows a different separation angle of 57 deg and MICH is bigger by factor of ~6.

  - REFL11 is dominated by PRC. The PRC signal is bigger than MICH by a factor of ~100, which agrees with the finesse simulation.

    However the separation angle between PRC and MICH are different. The measurement said only 19 deg, but the simulation said ~ 90 deg.

  - Woops, I forgot to calibrate the outputs from the LOCKIN module.

    The whole values must be off by a certain factor due to the lack of the calibration , but fortunately it doesn't change the demodulation phases.

Quote from #4884

I was able to measure the sensing matrix in the PRMI configuration.

The results will be posted later.

 

  4898   Tue Jun 28 14:21:41 2011 kiwamuUpdateIOOsomething wrong ? : Power incident on REFL11 and REFL55

The measured change in the REFL DC power with and without PRM aligned seems unacceptably small.  Something wrong ?

The difference in the power with and without PRM aligned should be more than a factor of 300.

         [difference in power] = [single bounce from PRM] / [two times of transmission through PRM ]

                                          = (1-T) / T^2 ~ 310,

where T is the transmissivity of PRM and T = 5.5% is assumed in the calculation.

Also the reflectivity of MICH is assumed to be 1 for simplicity.

Quote from #4894

We now have (with the PRM misaligned):

REFL11:  Power incident = 7.60 mW ;  DC out = 0.330 V  => efficiency = 0.87 A/W

REFL55:  Power incident = 23 mW ;  DC out = 0.850 V  => efficiency = 0.74 A/W

and with the PRM aligned::

REFL11:  DC out = 0.35 V  => 8 mW is incident

REFL55: DC out = 0.975 V  => 26 mW is incident

 

  4897   Tue Jun 28 11:25:56 2011 AlastairBureaucracyComputersrestarted elog

The manual instructions on the 40m wiki for restarting wouldn't work.  I killed the process okay, but then I got an error saying it "couldn't bind to port 8080, please try again using -p to select port".  The automated script worked though.

  4896   Tue Jun 28 10:11:13 2011 steveUpdateIOOPower incident on REFL11 and REFL55

Quote:

I measured the power incident on REFL11 and REFL55.  Steve was concerned that it is too high.  If we consider this elog the incident power levels were REFL11: 30 mW and REFL55: 87 mW. (assuming efficiency of ~ 0.8 A/W @1064nm for the C30642 PD).  However, currently there is a combination of Polarising BS and Half-waveplate with which we have attenuated the power incident on the REFL PDs.  We now have (with the PRM misaligned):

REFL11:  Power incident = 7.60 mW ;  DC out = 0.330 V  => efficiency = 0.87 A/W

REFL55:  Power incident = 23 mW ;  DC out = 0.850 V  => efficiency = 0.74 A/W

and with the PRM aligned::

REFL11:  DC out = 0.35 V  => 8 mW is incident

REFL55: DC out = 0.975 V  => 26 mW is incident

These power levels may go up further when everything is working well.

The max rated photo-current is 100mA => max power 125mW @0.8 A/W.

 

What is the power level on MC_REFL_ PDs and WFS  when the MC is not locked?

  4895   Tue Jun 28 09:50:03 2011 SureshUpdateIOOWFS2 RF response

The WFS2  Transimpedance has been measured to determine if it also suffers from the same 200MHz oscillations seen in WFS1 sensor head

The attached plots (pdf attached) show that the 29.5 MHz peak needs tweaking in Q2 and Q1 seems to have a much lower transimpedance than other quadrants.  The table below summarises the resonances and notches of the ckt

 

WFS2 Pole(MHz) Z(Ohms) Notch (MHz) Z(Ohms)
Q1 29.63 576 59.25 0.76
Q2 28.5 862 59.0 1.41
Q3 29.8 766 59.25 1.05
Q4 29.8 704 59.0 1.09

 

The peak at 10MHz is much sharper than the similar peak at 13MHz in the case of WFS1.  Is this a matter for some concern? 

The 200MHz oscillation once again exists in Q2, Q3 and Q4.  This sensor head will also require the same treatment as WFS1.

 

 

 

Quote:

I have shifted the Jenne laser back to the small table where we do RF PD characterisation (RFPD table).  I found several 25pin D-type connector cables, connected them in tandem and am using that to power the WFS2 sensor head at the RFPD table. 

The set up is ready for looking at the RF response of the  WFS sensors.  Will continue tonight.

 

 

Attachment 1: WFS2_transimpedance.pdf
WFS2_transimpedance.pdf WFS2_transimpedance.pdf WFS2_transimpedance.pdf WFS2_transimpedance.pdf
  4894   Tue Jun 28 07:46:54 2011 SureshUpdateIOOPower incident on REFL11 and REFL55

I measured the power incident on REFL11 and REFL55.  Steve was concerned that it is too high.  If we consider this elog the incident power levels were REFL11: 30 mW and REFL55: 87 mW. (assuming efficiency of ~ 0.8 A/W @1064nm for the C30642 PD).  However, currently there is a combination of Polarising BS and Half-waveplate with which we have attenuated the power incident on the REFL PDs.  We now have (with the PRM misaligned):

REFL11:  Power incident = 7.60 mW ;  DC out = 0.330 V  => efficiency = 0.87 A/W

REFL55:  Power incident = 23 mW ;  DC out = 0.850 V  => efficiency = 0.74 A/W

and with the PRM aligned::

REFL11:  DC out = 0.35 V  => 8 mW is incident

REFL55: DC out = 0.975 V  => 26 mW is incident

These power levels may go up further when everything is working well.

The max rated photo-current is 100mA => max power 125mW @0.8 A/W.

 

  4893   Tue Jun 28 02:11:47 2011 JenneUpdateLockingLatest MICH noise budget

I have measured / calculated the latest MICH noise budget.  It doesn't really look all that stellar.

MICH_noise_budget_as_of_28June2011.png

As you can see, we are nowhere near being shot noise limited, since there's a huge discrepancy between all of the measured spectra and the teal Shot Noise line. 

One possible suspect is that the analog whitening filters weren't on when I took my measurements.  I didn't actually check to ensure that they were on, so they might not have been.  Right now we're limited by electronics and other boring noises, so I need to make sure we're limited by the noise of the diode itself (we don't have enough light in the IFO to actually be shot noise limited since that takes 2.5mA for AS55 and I only have 1.1mA, but we should be ~within a factor of 2ish).

  4892   Tue Jun 28 01:18:53 2011 ranaHowToSUSfree swing all optics

Chris Wipf tells me that the EPICS Mutex Jumbo Mumbo can be overcome by upgrading our EPICS. We should get one of Jamie's assistants to get this going on one of the Ubuntu workstations.

  4891   Mon Jun 27 16:57:06 2011 steveUpdatePhotosHaixing is back

He has moved the levitation stuff for his surf student to Jan's lab in W-Bridge.

Attachment 1: P1070914.JPG
P1070914.JPG
  4890   Mon Jun 27 10:04:29 2011 kiwamuSummaryGeneralWeekly report

 Summary for the week ending June 26th.  (Number of elog entries = 53)

- SUS
  A BIO installed on 1X2.
  A peak finding script was prepared for diagonalization of the OSEM input matrices
  The suspension readout coefficients were changed to have unit of [um] and [urad] in each signal.
 
- ABSL
    LWE NPRO controller was brought by Peter King.
    The I-P cuvre and beam profile was measured. Nominal current was chosen to 1.8 [A].
    The access tube between PSL and AS table was back in place.
 
-RFPD
   The REFL55 characterization was analyzed (impedance gain = 615 Ohm, shot noise intercept current = 1.59 mA )
 
- MC
   WFS1 check, the 29MHz resonance need to be adjusted.
   The MC locking gain was increased by 6 dB to avoid an oscillation at 30 kHz.
 
- LSC
  The sensing matrices were measured in DRMI configuration and PRMI configuration
 
- Fiber experiment
   QPDY_PD was repositioned to accommodate the fiber stuff on the ETMY table.
   Succeeded in introducing the IR beam into the fiber coupler.
 
- TT characterization
    Th optic bench next to MC2 was cleaned up and leveled
 
- Vent list wiki page
   A wiki page was made for the vent detailed plan.
 
- CDS
  A foton's malfunction was found. It can run correctly only on Pianosa.
  Some Dell machines were gone to Rod Luna
 
- 40m specfic safety training for the SURFs
  4889   Mon Jun 27 00:23:11 2011 ranaUpdateCDSETMX SIDE problem

The slow readback of the ETMX side seems to also have something flaky and bi-stable. This is not an issue for damping, but it disables the SIDE watchdog for ETMX and makes it unsafe if we accidentally use the wrong damping sign.

Attachment 1: etmx-side.png
etmx-side.png
  4888   Sun Jun 26 22:38:20 2011 ranaUpdateCDSMC1 LR dead for > 1 month; now revived temporarily

 Since the MC1 LRSEN channel is not wasn't working, my input matrix diagonalization hasn't worked today wasn't working. So I decided to fix it somehow.

I went to the rack and traced the signal: first at the LEMO monitor on the whitening card, secondly at the 4-pin LEMO cable which goes into the AA chassis.

The signal existed at the input to the AA chassis but not in the screen. So I pressed the jumper wire (used to be AA filter) down for the channel corresponding to the MC1 LRSEN channel.

It now has come back and looks like the other sensors. As you can see from this plot and Joe's entry from a couple weeks ago, this channel has been dead since May 17th.

The ELOG reveals that Kiwamu caught Steve doing some (un-elogged) fooling around there. Burnt Toast -> Steve.

bt.jpg

993190663   =      free swinging ringdown restarted again

Attachment 1: lrsen.png
lrsen.png
  4887   Sun Jun 26 18:35:16 2011 ranaHowToSUSfree swing all optics

I used scripts/SUS/freeswing-all.csh to give the optics a kick and then turn off their watchdogs and collect the free swinging data.  Final script end time = 993173551. Start taking data ~ 993173751

I had to fix up the script a little: it had amateur stuff in there, such as undefined variables.

It still doesn't work that well. On the new Ubuntu workstations, pianosa, it fails by just not setting some of the EPICS variables using the EZCA stuff.

On Allegra, it failed on ~1 out of 10 commands by returning "epicsThreadOnce0sd epicsMutexLock failed" ???

On Pianosa, it sometimes says, instead, "epicsThreadOnceOsd: pthread_mutex_lock returned Invalid argument.".   Ah...now I understand?

So finally, I had to run the script on op340m to get it to actually run all of its commands. That's right; I used a 15 year old Solaris 9 Blade 150 because none of our fancy new Linux machines could do the job reliably.

Fixing our EZCA situation is a pretty high priority; if the locking scripts fail to run ~1 command every hour its going to completely derail the lock acquisition attempts.

If you want to use the IFO tonight, just run the script again on op340m again when you're done.

Attachment 1: ringdown.png
ringdown.png
  4886   Sun Jun 26 16:17:22 2011 ranaUpdateCDSdiagonalization of MC input matrix

I have updated the scripts/SUS/peakFit/ directory so that it now finds the SUS input matrix coefficients in addition to just finding the free-swinging peaks.

Procedure:

  1. Get OSEM sensors data via NDS2 from a time when the optics have been kicked and then left free swinging.
  2. Downsample the data to 64 Hz and save.
  3. Make power spectra with a 1 mHz resolution (i.e. we need a few hours of data) and  ~10 averages.
  4. use the fminsearch lorentzian peak fitter -> save the peak frequencies
  5. Make Transfer Function estimate matrix at the peak frequencies between all OSEMs (this makes a 5x4 complex matrix)
  6. The matrix should be real, so make sure its mostly real and then take the real part only
  7. Normalize (height of biggest peak for each f_DOF should be 1)
  8. Add a Butterfly mode vector. This makes the sensing matrix go from 5x4 to 5x5. (Butterfly a.k.a. Pringle)
  9. Invert
  10. Normalize so that the biggest element in each Sensor2DOF column is 1.
  11. Load values into MEDM screen and then verify by another free swinging data run.

 

The attached PDF shows how much rejection of the unwanted DOFs we get between the existing diagonal input matrix and this new empirical matrix. Previously, the decoupling was only a factor of a few for some of the modes. Now the decoupling is more like orders of magnitude (at least according to this calculation). It will be worse when we load it and then try another free swinging run. However, the fact that the suppression can be this good means that the variation in the coefficients at the ~hours time scale is at least this small (~< 0.1%)

 

That's the basic procedure, but there are a lot of important but mainly technical details:

  1. Free swinging data must be taken with the angle bias ON. Otherwise, we are not measuring the correct sensing gain (i.e. the magnets are not in their nominal place within the OSEM-LED beam)
  2. Data must be checked so that the shadow sensor outputs are in their linear regime: if they are exploring the cubic part, then the fundamental is being suppressed.
  3. Instead of just using the peak frequency, I average a few points around the peak to get better SNR before inversion. I think this will make the results more stable.
  4. All previous input matrix diagonalization efforts (Buckley, Sakata & Kawamura, Black, Barton, Gonzalez, Adhikari & Lawrence, Saulson,...) for the past ~15 years have been using the spectra's peak height data. Today's technique uses the TF and so is more precise. The coherent transfer function is always better than just using the magnitude data.
  5. This method is now fairly automatic - there's no human intervention in fudging values, choosing peak heights, frequencies, etc.
  6. We'll have to rerun this, of course, after the mirrors are aligned and after the OSEM whitening fiasco is cleaned up somewhat.

I'll set the optics to be aligned and then swing tonight.

Attachment 1: inMatDiag.pdf
inMatDiag.pdf
  4885   Sun Jun 26 16:02:12 2011 kiwamuUpdateIOOFriday MC activity

[Rana / Kiwamu]

 Last Friday we did several things for MC :

   - aligned the incident beam to MC

   - increased the locking gain by 6 dB and modified the auto-locker script accordingly

   - improved the alignment of the beam on the MC_REFLPD photo diode

 


(Motivation)

 In the beginning of the work, we wanted to know what RF frequency components are prominent in the reflection from MC.

Since the WFS circuits are capable for two RF notches, we wanted to determine which frequencies are appropriate for those notches.

So for the purpose we tried searching for unwanted RF components in the reflection.

However during the work, we found several things that needed to be fixed, so we spent most of the time for improving the MC locking.

 

(Some notes)

 - Alignments of the incident beam

At the beginning, the reflection from MC was about 2.2 in C1:IOO-REFLDC and the lock of MC had been frequently unlocked.

This situation of high reflection seemed to be related to a work done by Suresh (#4880).

Rana went to the PSL table and tweaked two input steering mirrors in the zig-zag path, and finally the reflection went down to ~ 0.8 in C1:IOO-REFLDC.

This work made the lock more robust.

 

 - Change of the locking gain

 After the alignment of the incident beam, we started looking at the time series of the MC_REFLPD signal with an oscilloscope as a start point.

What we found was a significant amount of 30 kHz components. This 30 kHz oscillation was thought be a loop oscillation, and indeed it was so.

We increased the loop gain by 6 dB and then the 30 kHz components disappeared successfully.

So the nominal locking gain of MC is now 11 dB in C1:IOO-MC_REFL_GAIN. The auto locker script was also modified accordingly.

 

- RF components in the MCREFL signal

After those improvements mentioned above, we started looking at the spectrum of the MCREFL PD using the spectrum analyzer HP8590.

The 29.5 MHz component was the biggest components in the spectrum. Ideally this 29.5 MHz signal should be zero when MC is locked.

One possible reason for this big 29.5 MHz signal was because the lock point was off from the resonant point.

We tweaked the offset in the MC lock path using a digital offset, C1:IOO-MC-REFL_OFFSET.

We found an offset point where the 29.5MHz signal went to the minimum, but didn't go to zero.

 

(works to be done)

So it needs some more works to investigate the cause of nonzero 29.5 MHz signal as well as investigation of what RF components should be notched out.

A good start point would be writing a GPIB interface script such that we can get the spectra from HP8590 without any pains.

  4884   Sat Jun 25 06:09:38 2011 kiwamuUpdateLSCFriday locking

I was able to measure the sensing matrix in the PRMI configuration.

The results will be posted later.

  4883   Sat Jun 25 04:40:43 2011 SureshUpdateASCWFS1 Transimpedance measurement

WFS1 Transimpedance

The attached plots show the location of the ~29.5 MHz pole and the 59 MHz notch for each quadrant of the WFS1 Sensor head.

 

WFS1 Pole (MHz) Z(Ohms) Notch (MHz) Z(Ohms)
Q1 28.89 598 60.38 0.83
Q2 29.20 513 57.70 0.57
Q3 29.63 681 59.63 0.89
Q4 28.89 609 58.13 0.78

 

As may be seen from the above table, these frequencies will need to be adjusted in some cases.

From the plots we can see that, when there is no attenuation set on the attenuator AT65-0263 (ref D990249-A),  the MAX4107  oscillations are seen in Q2,Q3,Q4 quadrants at around 200 MHz.  

Rana suggested, from his previous encounter with this circuit, that the solution is to remove the second MAX4106 and the attenuator on the RF line to avoid this oscillation.

 

WFS1_transimpedance.pdf

A look at the circuit board shows that some of the inductors have not been mounted.  That explains the presence of only one notch though the schematic shows two. 

P6250224.JPG         P6250223.JPG

  4882   Sat Jun 25 00:00:28 2011 SonaliUpdateGreen LockingFibre Coupling.

 What I did today.

1. I tried to align the IR input beam by aligning the two mirrors, to couple input light into the fibre.

2.I was unsuccessful for a long time even though I tried a lot of tricks.

3. I also tried to use the optical fault locator to superpose the IR beam spot onto the beam spot of the other laser to facilitate effective coupling.

4.But the crucial point was to superpose the input beam path in the perfect direction of the output beam path and not just the beam spot.(the input cone and the output cone are perfectly aligned).

5.After one whole day of trial and thought, I managed to couple light into the fibre, and saw the output beam spot on the screen-camera-monitor set-up which we had arranged. Eurekka !!;)

6.I then used a power meter to measure the input beam power and the output beam power.

7.It was a disappointing 2% . I had read in project reports of many students of a 20% success.

8.After a lot of subtle tweaking of the mirrors using the knobs, I managed to increase the percentage of output beam to 12%.

9. This is a workable level.

10.A day of lot of new learning! Pictures of the setup are attached.:)

 

Attachment 1: Fibre_coupling_successful_24_june.jpg
Fibre_coupling_successful_24_june.jpg
Attachment 2: Beam_output_on_screen.jpg
Beam_output_on_screen.jpg
  4881   Fri Jun 24 22:35:23 2011 ranaConfigurationCDSdataviewer broken on pianosa

When I try to get minute trend, it says "word too long".

  4880   Fri Jun 24 21:21:46 2011 SureshUpdatePSLPlaced labels on the zig-zag mirrors on PSL table

I put labels on the pair of beam steering mirrors which are at the output end of the PSL table.  I had changed one of these mirrors (elog) and Jenne had changed the other (elog).  This was at about 3PM today

I just learned from Kiwamu that this has messed up the MC alignment.

 

  4879   Fri Jun 24 17:04:25 2011 NicoleUpdateSUSBasic Laser Safety Training; Moved TT Mirror; Horizontal Displacement Mech Plan

Today Ishwita, Sonali, and I completed basic laser safety training with Peter King. I completed the Laser Safety Quiz and have turned in my certificate sheet.

I just need to turn in a signed copy of the Lab Safety Checklist to SFP (which I can now have signed by Koji after completing the course).

 

Steve and I have removed the TT mirror from the clean box. It is now on the small optical table in the lab that I have been working on.  Thanks to Steve, all of the mechanical components for the horizontal displacement measurement experiment are compiled and on the small optical table. Here is a photo of the small optical table with the gathered components. CompiledParts.JPG

The plan is to attach the slider and the shaker directly to the black mounting plate. On the slider, we we then place the smaller black mounting plate (with the lip). The lip will attach to the shaker. We know exactly where to drill and everything is lined up. The shaker will be placed on the smaller black mounting plate (with the lip).  The assembly will begin on Monday.

 

Here is a photo of the planned set-up for the shaker and the horizontal slider + mounting base.

 HorizontalDispMount.JPG

  4878   Fri Jun 24 10:38:01 2011 steveUpdateCamerasITMY camera gets fixed

ITMY gets new Tamron M118FM50 that has improved close focusing. It is a small fixed focal length camera so the video tube cover can be put on.

The Watec LCL-902K 1/2" ccd camera was losing it power supply voltage because of bad connection. It was replaced.

  4877   Fri Jun 24 07:49:23 2011 steveUpdateABSLI-P curve of LWE with serial numbers

Quote:

 The I-P curve was measured again, but this time in a lower current range of 1.0-1.9 [A].

Quote from #4842

The old days the NPRO ( inside the MOPA ) was running ~1.7A  500 mW

 Lightwave Laser Head M126-1064-700   sn238,  mounted on full size Al base and side heat sink on

Controller 125/126 Smart Supply   sn 201M

  4876   Fri Jun 24 07:40:23 2011 steveUpdateSUSPRM damping restored

The PRM sus damping restored. C1:SUS-PRM_SDPD_VAR is still 20-30mV and going up.  Side gain  turned on. This pulled it down to 5-8 mV

Why is the side osem sensing voltage 4.4V ? It can not be higher than ~2.4V.......something is rotten in the state of Denmark?

Edit by KI:

 It's because Valera increased the transimpedance gain of the PRM SIDE OSEM to match the signal level to the new ADC range (#3913 ).

  4875   Fri Jun 24 01:05:32 2011 NicoleSummarySUSTransfer Function Model Analysis Summary and New Posted LED V vs. I Curve

I have updated the TT suspension wiki to include a new page on my transfer function model. In this new page, an introduction and analysis of my transfer function (including a comparison of the transfer functions for a flexibly- and rigidly-supported damper) are included.  This page contains linear and logarithmic bode plots.  Here is a link to the transfer function page.

 

I have also updated my photosensor page on the TT suspension wiki so that the experimental data points in my current versus voltage plot are plotted against the curve provided by the Hamamtsu data sheet. I have also included an introduction and analysis for my mini-experiment with the forward voltage and forward current of the LED. Here is link to the photsosensor page.

  4874   Fri Jun 24 00:13:24 2011 kiwamuUpdateABSLbeam profile measurement of LWE

The beam profile of the LWE (LightWave Electronics) NPRO was measured.

Mode matching telescopes will be designed and setup soon based on the result of the measurements.

 

Here is a plot of the measured beam profile.

beam_profile.png

 (some notes)

The measurement was done by using Kevin's power attenuation technique (#3030).

An window was put just after the NPRO and the reflected beam was sampled for the measurement to avoid the beam scan saturated.

  4873   Thu Jun 23 23:54:29 2011 KojiOmnistructureEnvironmentWe are saved

Sonali, Ishwita, and another anonymous SURF saved the long-lasted water shortage of the 40m

Attachment 1: IMG_0023.jpg
IMG_0023.jpg
  4872   Thu Jun 23 22:59:45 2011 kiwamuUpdateABSLI-P curve of LWE

 The I-P curve was measured again, but this time in a lower current range of 1.0-1.9 [A].

The plot below is the latest I-P curve.

IP_curve_small.png

(Decision)

Based on the measurement and some thoughts, I decided to run this laser at about 1.8 [A] which gives us a middle power of ~ 360 [mW].

In the 40m history, the laser had been driven at 2.4 [A] in years of approximately 2006-2009, so it's possible to run it at such a high power,

but on the other hand Steve suggested to run it with a smaller power such that the laser power doesn't degrade so fast.


(notes)

  The laser controller handed from PK (#4855) was used in this measurement.

The nominal current was tuned to be 1.8 [A] by tuning a potentiometer on the laser head (see page.18 on the manual of LWE).

There was a huge bump around 1.4 [A] and sudden power drop at 1.48 [A] although I don't know the reason.

Quote from #4842

The old days the NPRO ( inside the MOPA ) was running ~1.7A  500 mW

  4871   Thu Jun 23 22:53:02 2011 kiwamuUpdateCDSran activateDQ.py

I found some DQ channels (e.g. SENSOE_UL and etc.) for C1SUS haven't been activated, so I ran activateDQ.py.

Then I restarted daqd on fb as usual. So far the DQ channels look working fine.

  4870   Thu Jun 23 22:39:34 2011 JenneUpdatePSLPMC found unlocked

I found the PMC unlocked.  Koji noticed that the FSS Slow Actuator Adjust was railed at the positive end of the slider.  I set it close to zero, and relocked the PMC.  The FSS slow loop servo is doing its thing, and the PMC and MC are now locked. 

  4869   Thu Jun 23 22:00:22 2011 JamieUpdateSUSburt snapshot

I recorded a burt snapshot of these settings: /opt/rtcds/caltech/c1/burt/autoburt/snapshots/2011/Jun/23/21:40

  4868   Thu Jun 23 21:35:46 2011 Jamie, Rana, KiwamuUpdateSUSFix calibration for sus sensors

We have fixed the counts-to-micron (cts2um) calibration for the suspension sensor filters. Each suspension sensor filter bank (e.g. ULSEN) has a "cts2um" calibration filter. These have now been set with the following flat gains:

   40 V       10^3 um         um
 -------- *  --------  = .36  --
 2^16 cts     1.7 V           ct

The INMTRX was also fixed with proper element values:

UL UR LR LL SIDE  
.25 .25 .25 .25 0 POS
1.666 1.666 -1.666

-1.666

0 PIT
1.666 -1.666 -1.666 1.666 0 YAW
0 0 0 0 1 SIDE

This was done for all core optic suspensions (BS, PRM, SRM, ITMX, ITMY, ETMX, ETMY).

 

  4867   Thu Jun 23 21:34:21 2011 kiwamuUpdateCDSno foton on the CentOS machines

For some reasons foton's deafault sample rate is NOT correct when it runs on the CentOS machines.

It tries to setup the sample rate to be 2048 Hz instead of 16384 Hz until you specify the frequency.

To avoid an accidental change of the sample rate,

running foton on CentOS is forbidden until any further notifications.

Run foton only on Pianosa.

 

Additionally I added an alias sentence in cshrc.40m such that people can not run foron on CentOS (csh and tcsh, technically speaking).

Below is an example of raw output when I typed foron on a CentOs machine.

    rossa:caltech>foton
    DO NOT use foton on CentOS

  4866   Thu Jun 23 10:35:12 2011 steveUpdateComputersold computers leaving the lab

Rod Luna picked up these computers for Larry Wallace yesterday: Dell Inspiron 530, Dell Dimension 4600 and SunBlade 1000

Attachment 1: P1070905.JPG
P1070905.JPG
Attachment 2: P1070904.JPG
P1070904.JPG
Attachment 3: blade1000.JPG
blade1000.JPG
  4865   Thu Jun 23 10:17:49 2011 steveUpdatePSLAP table is open to PSL again

Access to the north side of the PSL table is blocked by the    8" beam guard. This opens the beam pathways between them.

Attachment 1: P1070911.JPG
P1070911.JPG
Attachment 2: P1070910.JPG
P1070910.JPG
  4864   Thu Jun 23 09:46:16 2011 ranaUpdateLSCPRMI locking : not stable enough

All the suspensions are bad until you fix them. But, ... there is a script which can be used to diagnose them today:

Python SUStest

  4863   Thu Jun 23 05:22:51 2011 kiwamuUpdateLSCPRMI locking : not stable enough

I was trying to measure the sensing matrix in the PRMI configuration, but basically gave up.

It is mainly because the lock of PRMI wasn't so stable and it didn't stay locked for more than a minute.

It looked like an angular motion fluctuated a lot around 1- 3 Hz. The beam spot on the AS camera moved a lot during the lock.

I have to figure out who is the bad suspension and why.

  4862   Thu Jun 23 02:12:12 2011 SonaliUpdateGreen Lockingwork schedule.

 June 22-June 24:

1.Coupling light into fibre at the ETMY.

2.Routing of the fibre to the PSL table.

June 27-June 30:

1.PSL optical table layout sketching.

2.Combining the PSL beam with fibre output onto a BS and then superpose them on a New Focus 1611 PD.

July 5-July 8:

1.Conversion of the PD output to voltage using MFD(Mixer Frequency Discriminator).

2. Report writing.

July 7: 5:00 pm: 1st Report Due.

July 11-July 22:

1.Locking Y-arm to PSL.

2.Setting up the feedback loop using the MFD output as the error signal and acting on the AUX laser frequency.

July 25-Aug 5:

1.Y-Arm cavity characterisation.

Measurement of the transmission of IR and green light through the cavity.

2.Analysis.

To obtain FSR, Finesse,Loss of the Cavity, Visibility, Transverse Modes(g-factor, astigmatism), Reflectivity, Q-factor.

3.Report and abstract writing.

Aug 1: 5:00 pm: 2nd Report and absract due.

Aug 8-11:

Preparation for talk and seminar.

  4861   Wed Jun 22 21:36:41 2011 ranaSummaryGeneralJuly 2011 vent plan

I put a paper Peet's bag with half of the Mini-Moos into George.

  4860   Wed Jun 22 18:51:47 2011 SureshUpdateIOOWFS2 RF response

I have shifted the Jenne laser back to the small table where we do RF PD characterisation (RFPD table).  I found several 25pin D-type connector cables, connected them in tandem and am using that to power the WFS2 sensor head at the RFPD table. 

The set up is ready for looking at the RF response of the  WFS sensors.  Will continue tonight.

 

ELOG V3.1.3-